OpenStudy (anonymous):

Find the Domain of the function f(x)=5/x+2

OpenStudy (precal):

x can not equal 0 f(x)=(5/x)+2 but if it is f(x)=5/(x+2) then x can not be -2

OpenStudy (anonymous):

its f(x)=5/(x+2)

OpenStudy (precal):

ok x can not be -2

OpenStudy (anonymous):